i dont get pre-algebra that very good so can you help me with this problem:
4x-7=2x+15

Answer by josmiceli(19441)

4x-7=2x+15
Think of the equals sign as a balence point. You're balencing the left side with
the right side. like this:
stuff = stuff
junk = junk
junk + stuff = junk + stuff
also
junk + stuff = stuff + junk (you don't care about the order)
3x(junk) = 3x(junk) (it still balences)
so,
4x-7=2x+15 must be like stuff = stuff, right?
You balence this by taking away the same amount of stuff from both sides
or adding the same amount of stuff to both sides
first add 7 to both sides
4x -7 +7 = 2x +15 +7
now just do the addition
4x = 2x + 22
Now subtract 2x from both sides
4x - 2x = 2x - 2x + 22
do the subtraction
2x = 22
Now divide both sides by 2 (that doesn't change the balence either)
x = 11
that's the answer
You can check it by plugging it into the equation
you started with
4*(11) - 7 = 2*(11) + 15
do the multiplications
44 - 7 = 22 + 15
Is this true?
37 = 37
Yes, it is!
It's simple if you know the rules and you look at it
the right way. stuff = stuff